OpenStudy (zaarin):

On a blueprint, 1/4 inch represents 1 foot. If a window is supposed to be 60 inches wide, how wide would its representation be on the blueprint? A.7/6 inches B.5/4 inches C.4/3 inches D.15 inches E.56/3 inches

OpenStudy (***[isuru]***):

hi, I think u will get through this if u have the conversion factors 1 foot = 12 inches so now u can find the length given as 60 inch in foot and solve the problem can u do that bro ..?

OpenStudy (zaarin):

Sorry bro I have already tried it but that didn't work.

OpenStudy (anonymous):

12 inches =1 feet 1 inch=1/12 feet 60 inches=60*1/12 =5 feet now again by unitary method you can find size of blueprint.

OpenStudy (zaarin):

ok